OpenStudy (anonymous):

Solve the equation. x^4-x^2=2x^2+4 x= Solve the equation. x^4-x^2=2x^2+4 x= @Mathematics

OpenStudy (anonymous):

\[x^{4}-x^{2}=2x^{2}+4\]\[x^{4}-3x^{2}-4=0\]\[(x^{2}-4)(x^{2}+1)=0\]\[x=1, -1, 2, -2\]

OpenStudy (anonymous):

Wow... Even I myself was shocked of the answer... lol

OpenStudy (anonymous):

thanks for your speedy response

OpenStudy (anonymous):

given, x^4-x^2=2x^2+4 or, x^4-3x^2-4=0 or, x^4-4x^2+x^2-4=0 or, x^2(x^2-4)+1(x^2-4)=0 or, (x^2-4)(x^2+1)=0 or, x=2,-2

OpenStudy (anonymous):

1, and -1 are not real roots dont forget.

OpenStudy (anonymous):

Well \[1^{2}=1\]and \[(-1)^{2}=1\] lol... Sorry

OpenStudy (anonymous):

roots are 2,-2,i,-i

OpenStudy (anonymous):

I just get it wrong sometimes :P